### v1

# Quantum Geometry: Revisiting electronic scales in quantum matter

Abstract
Electronic properties of solids are often well understood via the low energy dispersion of Bloch bands, motivating single band approximations in many metals and semiconductors. However, a closer look reveals new length and time scales introduced by quantum dipole fluctuations due to interband mixing, which are reflected in the momentum space textures of the electronic wavefunctions. This structure is usually referred to as quantum geometry. The scales introduced by geometry not only qualitatively modify the linear and nonlinear responses of a material but can also have a vital role in determining the many-body ground state. This review explores how quantum geometry impacts properties of materials and outlines recent experimental advances that have begun to explore quantum geometric effects in various condensed matter platforms. We discuss the separation of scales that can allow us to estimate the significance of quantum geometry in various response functions.
I Introduction
The quantum many-body problem involves an exponentially large Hilbert space, making a full treatment impossible for the vast number of electrons typically found in a quantum material. Much of the analytic progress in the field has been achieved through low-energy theories that focus on a subset of degrees of freedom within a smaller, more manageable projected Hilbert space. While identifying this subspace is a challenge for many quantum materials, the first step is to determine the relevant energy, length, and time scales.
In a crystalline solid, the first relevant scale is the atomic lattice constant $a$, which allows for the definition of crystal momentum ${\bm{k}}$ that labels both band energies $\varepsilon({\bm{k}})$ and wavefunctions ${\psi({\bm{k}})}$. Much of the success in describing metals and semiconductors lies in the fact that electron transport can be described semiclassically, with the long wavelength behavior well captured by the band dispersion and its momentum derivatives such as the band velocity $v_{\bm{k}}\!\sim\!\partial_{\bm{k}}\varepsilon({\bm{k}})$ or band effective mass $1/m^{*}\!\sim\!\partial_{\bm{k}}^{2}\varepsilon({\bm{k}})$, combined with a phenomenological time scale which reflects how fast momentum relaxes in the presence of impurities or electron-electron interactions, $\tau$. Together with the Fermi velocity, it defines a mean free path of electrons $\ell_{\rm mr}=v_{F}\tau$, usually much larger than the lattice constant scale $a$.
In conventional condensed matter wisdom, the structure of materials at scales comparable with $a$ is usually considered unimportant to its collective behavior, justifying long-wavelength approximations such as effective field theories. Alternative lattice model approaches are simplified by taking a single site per unit cell, e.g., the Hubbard model, also reducing to a single band approximation. However, new insights from topology and quantum geometry say that, in fact, orbital mixing can lead to surprising qualitative changes in the physical behavior of materials. In this overview, we emphasize that orbital mixing across unit cells is the source of a new length scale that comes not from the band energies but by how electron wavefunctions change with momentum ${\partial_{\bm{k}}\psi({\bm{k}})}$. We refer to these quantities broadly as the quantum geometry of solids.
Quantum geometry in materials has a simple interpretation as the ground state dipole fluctuations [1]. This generalization to infinitely large systems follows from the well-understood case of dipole fluctuations in isolated atoms due to virtual level transitions [2]. Generally, these fluctuations can be used to characterize the size, shape, and angular momentum of atomic orbitals; however, in the lattice, qualitatively new orbitals can emerge from the quantum interference of distinct neighboring sites and are substantially richer than the isolated atomic problem. We argue that quantum geometry, with origin in virtual inter-band transitions can be used in solids to quantitatively describe the response of bound electrons in infinite lattices, even in metals where only part of the electrons are bound, while others are itinerant, plane-wave-like carriers. Strikingly, a quantum geometric description of bound electrons can be straightforwardly applied to systems without translation symmetry and beyond the single particle approximation. Therefore, it is a universal feature that systematically tracks the scale of dipole fluctuations.
As we overview here, dipole fluctuations have dramatic effects in linear and nonlinear responses of common materials, particularly those in low spatial dimensions such as graphene, transition metal dichalcogenides, or moiré heterostructures. Recent developments have also established that quantum geometry can play a fundamental role in the competition between correlated states, where the quantum geometry of the normal state can help stabilize exotic quantum orders such as fractionalized topological phases and superconductivity. In the following, we explain the concept of band geometry, focusing on dipolar fluctuations and their immediate consequences for the physical properties of quantum matter. We further discuss avenues to access the geometric tensor directly via different experimental probes and some of the intrinsic difficulties in their measurement.
Figure 1: Dipole fluctuations lead to uncertainty in electron position. The typical spatial spread of localized electrons is characterized by a length scale $\ell_{g}$ determined by the quantum geometry of the ground state. a. The hydrogen atom, where the ground state wavefunction is an $s$ orbital with spread $\ell_{a}^{2}=3a_{B}^{2}/4$ where $a_{B}$ is the Bohr radius. b. Monolayer ${\rm MoTe}_{2}$, a semiconductor with $\sim 1\rm eV$ band gap. The localized electrons are shared between neighboring sites with a large spread spanning multiple unit cells due to nontrivial orbital interference.
II What is Quantum Geometry?
Geometric phases lead to interference effects when states adiabatically traverse closed loops in parameter space. From the polarization of light [3], threading of a magnetic flux [4], or changes in atomic positions of molecules [5], geometric phases were understood to be general in quantum systems [6], arising from the Riemannian structure of the projected manifold of occupied quantum states. Provost and Vallee [7] introduced the quantum geometric tensor (QGT) to describe the adiabatic evolution in parameter space, showing that it is characterized not only by a phase rotation but also by a loss of the projected norm of the quantum states. Both effects are simultaneously characterized by the real and imaginary parts of the QGT. While Berry phase effects in electronic structure theory are now very well understood [8], the effects of the real part, known as the quantum metric, has remained mostly unexplored. Nonetheless, just as its imaginary counterpart, the real part is bound to play a significant role in quantum states of matter.
Fidelity susceptibility.
The definition of the QGT is fairly universal and has only two requisites: a parameter space $\{\lambda_{\mu}\}$ and a projector $\hat{P}$, defining a subset of quantum states in the Hilbert space,
$\hat{\mathcal{Q}}_{\mu\nu}=\hat{P}\partial_{\mu}\hat{P}\partial_{\nu}\hat{P}$ (1)
where and $\partial_{\mu}\equiv\partial/\partial_{\lambda_{\mu}}$. We use the symbol without a hat $\mathcal{Q}_{\mu\nu}\!=\!\mathop{\mathrm{tr}}[\hat{\mathcal{Q}}_{\mu\nu}]$ to indicate the trace over the internal states of $\hat{P}$. Consider first the projector into a single state of an Hamiltonian $\mathcal{H}$, $\hat{P}=|\psi\rangle\langle\psi|$. When perturbations $\mathcal{H}+\sum_{\mu}\lambda_{\mu}\mathcal{H}_{\mu}^{\prime}$ are introduced, the parameters $\lambda_{\mu}$ define a parameter space and Eq. (1) its geometric tensor. In this case, the QGT captures the sensitivity of the state $\left|\psi\right>$ to perturbations, or fidelity susceptibility [9].
The fundamental reason behind quantum geometry is that the wavefunction overlap at adjacent points in parameter space deviates from unity, even when the wavefunctions are normalized at each point. Berry [6] famously demonstrated that wavefunctions acquire infinitesimal phases during this evolution, $\langle\psi(d\bm{\lambda})|\psi\rangle\approx 1-id\lambda_{\mu}\langle\psi|\partial_{\mu}\psi\rangle$, that accumulate to a gauge invariant geometric phase when the path is closed. The phase is captured by the Berry curvature $\Omega_{\mu\nu}=i\epsilon_{\mu\nu}\partial_{\mu}\langle\psi|\partial_{\nu}\psi\rangle$. A more subtle aspect is that the loss of norm is also gauge-invariant and measurable [10, 11]. Expanding the overlap to second order, we find
$|\langle\psi(d\bm{\lambda})|\psi\rangle|^{2}=1-g_{\mu\nu}d\lambda_{\mu}d\lambda_{\nu}+\mathcal{O}(d\lambda^{4}),$ (2)
where $g_{\mu\nu}$ can be seen as a metric that quantifies the distance between nearby quantum states. The QGT tensor proves invaluable in contexts where wavefunctions change abruptly, such as during quantum phase transitions [12], conical intersections of energy landscapes [5] or near singular band crossings in electric band structure [13], at which points the susceptibility Eq. (3) diverges [14, 15].
An elegant application of the geometric tensor for quantum materials was introduced by Walter Kohn [1], who studied a many-body ground state $\hat{P}_{0}=\left|\psi_{0}\right>\left<\psi_{0}\right|$ under twisted boundary conditions. The twist shifts the momentum in the spatial direction $\mu$ by the adiabatic parameter $\kappa_{\mu}$, $p_{\mu}\rightarrow p_{\mu}-\hbar\kappa_{\mu}$. Kohn argued that in the thermodynamic limit, the susceptibility Eq. (1) distinguishes the ground state of a metal from an insulator, diverging in metals while remaining finite for insulators [16]. This distinction based on the localization of wavefunctions bypasses the reference to a spectral gap, which can be ill-defined in disordered and many-body systems [1].
Dipole fluctuations
The momentum shift introduced by Kohn can be interpreted as a susceptibility towards dipole transitions between states in $\hat{P}$ and states in the complementary projector $1-\hat{P}$. Since position and momentum are conjugate variables, $\partial_{\mu}\hat{P}\equiv i[\hat{r}_{\mu},\hat{P}]$, we may express Eq. (1) as
$\mathcal{Q}_{\mu\nu}={\rm tr}\left[\hat{P}(i[\hat{r}_{\mu},\hat{P}])(i[\hat{r}_{\nu},\hat{P}])\right]=\langle\hat{r}_{\mu}(1-\hat{P})\hat{r}_{\nu}\rangle$ (3)
where we have defined $\langle\cdot\rangle\!\equiv\!\mathop{\mathrm{tr}}[\hat{P}\cdot]$ as the expectation value at $T=0$. As it contains the dipole transitions between the ground state and excited states, an insulator can have nonzero quantum geometry as long as the ground state is not a position eigenstate. The diagonal elements of the real part of $\mathcal{Q}_{\mu\nu}$, i. e. the quantum metric $g_{\mu\nu}={\rm Re~{}}\mathcal{Q}_{\mu\nu}$ introduces a length scale $\ell_{g}=\sqrt{\mathop{\mathrm{tr}}g}$ associated with the size of dipole fluctuations in the system along the $\mu$ spatial direction. This fundamental insight was first discussed by Resta with the introduction of the “localization tensor” [17] and soon after by Souza, Wilkens and Martin [18].
One does not need to look far to interpret the quantum metric as a measure of dipole fluctuations. A straightforward yet instructive example is an isolated Hydrogen atom subjected to an external electric field applied along $\hat{x}$. The perturbed Hamiltonian, $H^{\prime}=H-e\hat{x}E$, modifies the ground state wavefunction from $|\psi_{0}\rangle$ to $|\psi_{0}(E)\rangle$. The quantum metric is related to the fidelity via $g_{xx}=\partial|\langle\psi_{0}(E)|\psi_{0}\rangle|^{2}/\partial E^{2}$ as $E\rightarrow 0$. Using second-order perturbation theory, the metric becomes $g_{xx}=\sum_{m\neq 0}|\langle\psi_{0}|\hat{x}|\psi_{m}\rangle|^{2}=\langle(\hat{x}-\langle\hat{x}\rangle)^{2}\rangle_{0}$. This confirms that the metric measures the dipole fluctuations, which for the ground state of the hydrogen atom corresponds to the spatial spread of the $s$ atomic orbital, $\ell_{g}^{2}\!=\!(3/4)a_{B}^{2}$, with $a_{B}$ the Bohr radius. This example shows us that the size of the atomic orbital is a scale introduced by the geometric tensor in a single atom. Importantly, as long as there is time-reversal symmetry, the imaginary part of Eq. (3) vanishes exactly.
The advantage of Eq. (3) lies in its applicability to macroscopic systems where the behavior of electrons is qualitatively different from those in the Hydrogen atom. This is particularly important in bands with topological invariants where quantum geometry provides us with the language to capture the scale $\ell_{g}$ of dipole fluctuations of bound electrons, see Fig.1. The geometric length scale $\ell_{g}$ is usually comparable to the lattice constant $\ell_{g}\sim a$ and has dramatic signatures such as protected boundary modes [19] or quantization of Hall conductivity [20, 21]. For free electrons in a magnetic field, $\ell_{g}$ is nothing more than the magnetic length $\ell_{g}^{2}=\ell_{B}^{2}=\hbar/eB$, placing a similar formal ground to the quantum Hall effect in the continuum or in lattice models.
Here, it is our goal to emphasize that quantum geometry is indispensable in modeling materials with multiple bands - it presents a systematic way of including quantum corrections to the dipole operator beyond the conventional semi-classical approach. There has been increasing evidence that materials can exhibit drastically different properties solely due to differences in their quantum geometry [22]. This is particularly true when the band dispersion is significantly flattened by interference induce by, for example, a moiré potential. Therefore, the identification of observables capable of directly measuring or quantifying quantum geometry in real materials has emerged as an important challenge in the field [23].
III Quantum Geometry in Band Theory
Figure 2: Kohn vs. single band metric. a. The Kohn metric uses the projected $\hat{P}^{0}$ and quantifies dipole transitions between occupied and unoccupied states. b. Quantum metric for a single isolated band with index $m$ using the band resolved projector $\hat{P}^{m}$ that includes all dipole transitions between band $m$ and all other bands $n\neq m$. The infinitesimal distance $|\langle u_{m,k}|u_{m,k+dk}\rangle|^{2}$ can be written as $1-g^{m}_{\mu\nu}dk_{\mu}dk_{\nu}$ where $g^{m}_{\mu\nu}=\sum_{n\neq m}\langle u_{m,k}|\hat{r}_{\mu}|u_{n,k}\rangle\langle u_{n,k}|\hat{r}_{\nu}|u_{m,k}\rangle$.
The localization tensor, defined in Eq. (3), quantifies the dipole matrix elements between the ground state and the excited states. Even in the case of non-interacting electrons, this formalism is often impractical because of the vast Hilbert space. However, from the simple case of an isolated Hydrogen atom, we can gain valuable intuition about the effects of dipole fluctuations and derive a geometric framework to apply in systems with multiple electrons. This approach comes with its challenges, as the definition of a dipole operator in an infinite periodic system has been settled only a few decades ago with the modern theory of polarization [24]. Central to this resolution were the cell-periodic parts of the Bloch wavefunction.
Due to Bloch’s theorem, the electronic wavefunction in a periodic lattice potential can be decomposed into a plane-wave and a cell-periodic part, $|\psi_{n{\bm{k}}}\rangle=e^{i{\bm{k}}\cdot{\bm{R}}}|u_{n{\bm{k}}}\rangle$, where $n$ is the band index, ${\bm{k}}$ the crystal momentum and ${\bm{R}}$ a Bravais lattice vector. Being eigenstates of a Hermitian operator, the Bloch states are orthogonal. However, this orthogonality does extend to the cell-periodic function $|u_{n{\bm{k}}}\rangle$ which can have nonzero overlaps for different momenta. This introduces a notion of distance between two cell periodic states, giving rise to a Riemannian structure within the Brillouin Zone (BZ) by the adiabatic parameter ${\bm{k}}$ [25].
To be more precise, we note that the quantum geometric tensor is the dipole correlation function that requires a projector. A natural choice is the projector to a Bloch state resolved in both band index and crystal momentum $\hat{P}_{n{\bm{k}}}$. However, it is important to notice that the dipole matrix element $\hat{r}_{\mu}^{mn}({\bm{k}},{\bm{k}}^{\prime})=\langle u_{m{\bm{k}}}|\hat{r}_{\mu}|u_{n{\bm{k}}^{\prime}}\rangle$ contains both intra- and interband contributions [26], (cf. Fig. 2),
$\hat{r}_{\mu}^{mn}({\bm{k}},{\bm{k}}^{\prime})=-\delta_{{\bm{k}},{\bm{k}}^{\prime}}\left[\mathcal{A}_{\mu}({\bm{k}})\right]_{mn}+i\delta_{mn}\partial_{\mu}\delta_{{\bm{k}},{\bm{k}}^{\prime}}\ ,$ (4)
containing the matrix elements of the Berry connection $\left[\mathcal{A}_{\mu}({\bm{k}})\right]_{mn}=i\langle u_{m{\bm{k}}}|\partial_{\mu}u_{n{\bm{k}}}\rangle$, which are not by themselves gauge invariant. This does not pose an issue for QGT since the complementary projector in Eq. (3) removes the diagonal contributions of the position operator, making it gauge-independent. It can be compactly written in the band and momentum resolved basis as
$\mathcal{Q}_{\mu\nu}^{n}({\bm{k}})=\left\langle\partial_{\mu}u_{n{\bm{k}}}\right|1-\hat{P}_{n{\bm{k}}}\left|\partial_{\nu}u_{n{\bm{k}}}\right\rangle.$ (5)
From $\mathcal{Q}_{\mu\nu}^{n}({\bm{k}})$ one can similarly define a momentum resolved quantum metric $g^{n}_{\mu\nu}({\bm{k}})={\rm Re~{}}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]$ and Berry curvature $\Omega^{n}_{\mu\nu}({\bm{k}})=2{\rm Im~{}}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]$ in the BZ. These two control the overlap and geometric phase between two infinitesimally close cell-periodic states $|u_{n,{\bm{k}}}\rangle$ and $|u_{n,{\bm{k}}+d{\bm{k}}}\rangle$ respectively.
Berry curvature effects are well known in the semiclassical theory of transport in terms of the anomalous velocity [8]. Originally derived by Karplus and Luttinger through second-order perturbation theory [27], the anomalous velocity was later recognized as a consequence of wavefunction overlaps [8]. By explicitly deriving the equations of motion for a wave packet in the presence of an electric field, they demonstrated that the intrinsic anomalous velocity is precisely the Berry curvature [28, 29]. This paradigm shift has revolutionized our understanding of the Anomalous Hall effect, offering a fresh perspective distinct from conventional mechanisms based on scattering theory [30]. Furthermore, it has opened pathways to quantized Hall conductance by utilizing the quantization of the integral of Berry curvature over one band. This stems from the relation $\int_{\rm BZ}{\rm Im}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]=4\pi\mathcal{C}\epsilon_{\mu\nu}$ where $\int_{\rm BZ}\equiv\int d^{d}{\bm{k}}/(2\pi)^{d}$ and $\mathcal{C}$ is the Chern number that can only be an integer.
In contrast to the Berry curvature, the quantum metric has remained largely unexplored, primarily due to the lack of an intuitive understanding. Moreover, challenges inherent in its definition create additional conceptual hurdles that complicate its interpretation and practical application. The band resolved QGT Eq. (5) corresponds to dipole transitions between the band $n$ and all other bands $m\neq n$, filled or empty, at a given point in the BZ. It is thus qualitatively different from the metric introduced by Kohn, where the projector in Eq. (3) is the ground state projector $\hat{P}_{0}$, see Fig.2. For band electrons, we can define $\hat{P}_{0}$ into all occupied single-particle states as
$\hat{P}_{0}=\int_{\rm BZ}\sum_{n}\Theta(\mu-\varepsilon_{n{\bm{k}}})|u_{n{\bm{k}}}\rangle\langle u_{n{\bm{k}}}|$ (6)
with $\mu$ the chemical potential. In particular, in the case of multiple filled bands, the ground state geometric tensor cannot be obtained by summing $\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})$ over filled bands and momenta. This is because, unlike Berry curvature, the quantum metric is not band additive [31, 32].
The additive property can be explained via Wannier functions [33]. The integral of the quantum metric over the Brillouin zone (BZ) for a single band, $\int_{\rm BZ}{\rm Re}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]=g^{n}_{\mu\nu}$, sets a lower bound on the spatial spread of the Wannier function [34]. When additional bands are included in the construction of the Wannier function, it is reasonable to expect a more localized Wannier state with a reduced spatial spread. However, since the quantum metric is inherently a positive quantity, the sum of the metric of individual bands within a projected manifold cannot be less than the metric of the full manifold. Thus metric cannot be additive, $g_{\mu\nu}\neq\sum_{n\in{\rm occ.}}g^{n}_{\mu\nu}$.
Setting these subtleties aside, the quantum metric and Berry curvature have a close relationship that becomes evident in the case of Landau levels. As we know, the Chern number arises from the non-commutativity of the projected position operators, $\hat{X}=\hat{P}\hat{x}\hat{P}$ and $\hat{Y}=\hat{P}\hat{y}\hat{P}$, with $[\hat{X},\hat{Y}]=i\ell_{B}^{2}$, where $\ell_{B}$ is the magnetic length. The non-commutativity of these operators naturally implies that the variance $\langle\hat{X}^{2}+\hat{Y}^{2}\rangle$ is subject to a lower bound. Remarkably, the lowest Landau level satisfies the so-called trace condition [35], in which this variance saturates the bound. In simpler terms, the cyclotron orbits of an electron in a Landau level have the smallest spatial spread capable of supporting the necessary winding.
Figure 3: Separation of scales in the tight-binding approximation and the emergence of a geometric scale. Panel a. shows the band structure for a free particle in a honeycomb lattice potential. Panel b. shows the optical conductivity.
Although real materials deviate significantly from the idealized limit of Landau levels, the connection between topology and geometry remains true. Topological bands inherently require extended Wannier states, as the quantum metric is subject to a lower bound imposed by the topological index. Thus topology is a sufficient condition for non-trivial quantum geometry. This statement can be made more precise through an estimate of the quantum metric derived from the dielectric constant. Within the space of all insulators, it has been shown that the quantum metric in topological insulators is significantly larger than that in trivial insulators with comparable band gaps [36].
Importantly, topology is sufficient but not necessary for quantum geometry in real materials. Unlike the Berry curvature, the metric is not limited to systems that break time-reversal or inversion symmetries. Quantum geometry is ubiquitous to all known materials. The key challenge lies in developing robust methods to extract it and quantify its impact on the observed behavior.
IV Where to find quantum metric?
The experimental challenge in measuring the metric is two-fold. First, the probe must couple to the symmetric part of the QGT, and second, it must identify the specific geometric scale being probed. This issue ties back to the freedom in the choice of the projector $\hat{P}$ and the complementary projector $1-\hat{P}$ in the definition of the QGT in Eq. (3). Thus, the geometric tensor should not be viewed as an intrinsic property of a band or the ground state but rather as a property of the choice of projected states. These states are determined by the experimental probe, which itself imposes an effective truncation of the Hilbert space. That is, it is natural for the experimental probe to select the appropriate scale. A probe operating at some frequency range $\Delta\omega_{\rm ext}$ will carry an intrinsic time uncertainty $\Delta t\sim 1/\Delta\omega_{\rm ext}$. Alternatively, the Coulomb interaction strength $U$ can itself impose the energy window and which states can be truncated out of the Hilbert space.
This section is divided into three parts. The first part uses optical sum rules to establish the connection between response and geometry, and with this identify characteristic geometric length and energy scales. Building on this insight, we then review experiments capable of probing quantum geometry, with a clear identification of the associated projected subspace. Finally, focusing on a set of low-energy bands, we examine the effects of quantum geometry in many-body calculations.
IV.1 Instantaneous response and sum rules
The magnitude of dipole fluctuations in a real material depends on the time scale. In a metal, the dynamics of a single electron is diffusive in the long-time limit $t>\tau$ with $\langle r(t)r(0)\rangle\sim Dt$ with $D$ the Drude weight dictated by the band effective mass and carrier density. This diffusive propagation of fluctuations saturates at the mean free path, $\ell_{\rm mr}^{2}\sim D\tau$. Note for momentum preserving, clean metals $\ell_{\rm mr}\to\infty$. On the other hand, the very short-time dynamics $t<\hbar/V$ is governed by the local quantum chemistry of the atom. Quantum geometry introduces a length scale that lies in between these well-understood scales. It is comparable to the bandwidth of the optically active bands and is relevant for dynamics at times shorter than $\tau$ but larger than $\hbar/V$. We can formalize this intuition by rewriting the Souza-Wilkens-Martin (SWM) sum rule in terms of a Fermi surface (Drude) contribution $g_{\rm FS}$, a quantum geometric part due to sharing electrons across multiple orbitals $\ell_{g}^{2}$, and an atomic contribution $\ell_{a}^{2}$,
[…]
where $\sigma_{\mu\mu}(\omega)$ is the longitudinal optical conductivity along one spatial direction. We have introduced a normalization with respect to charge density $n$ to make a direct comparison of this scale of dipole fluctuations to the average inter-atomic distance $a\equiv n^{-1/d}$. This gives units of length squared to the integral at every dimension. The Fermi surface contribution, $g_{\rm FS}$, is either infinite or zero, depending on whether a Fermi surface is present or absent [37]. There is no such divergence for insulators, and the largest scale of dipole fluctuations is controlled by $\ell_{g}$. Fluctuations also exist at a smaller scale, $\ell_{a}$, corresponding to atomic orbitals. Capturing these requires integrating the optical conductivity up to a large energy cutoff $\Omega$ approximately given by the atomic level gap, to remove the contributions of core electron excitations. These high-energy transitions and the corresponding $\ell_{a}\sim a_{B}$ fluctuations are neglected in tight-binding approximations.
At intermediate frequencies between $\omega\in[\tau^{-1},\Omega]$, we can identify the geometric contribution to the localization tensor of the ground state $\ell_{g}$. This contribution is dominated by the band geometry without taking into account filling. The spectral weight at these frequencies would be pushed smoothly into the Drude peak and disappear if we adiabatically take electrons apart into the atomic limit, which means that it appears from the quantum interference and delocalization of orbitals across lattice sites. It can be viewed as the size of “molecular orbitals”. Importantly,the inter-band optical absorption yields a finite value of $\ell_{g}$, both in insulators and in metals. The emergence of this scale can be seen in Fig. 3, where we show the optical absorption for a free particle subjected to a honeycomb lattice potential. The finite energy contribution to optical conductivity, centered at $\omega_{g}$, arises from the quantum interference between the two sublattices, implying that $\ell_{g}$ is comparable to the atomic distances, and the energy $\omega_{g}$ comparable to the hopping energy between sublattices. Applying Eq. (7) to free electrons under a magnetic field of magnitude $B$, we find that $\omega_{g}=\omega_{c}=\hbar/m\ell_{B}^{2}$ and $\ell^{2}_{g}=\ell^{2}_{B}=\hbar/eB$.
Sum rules reveal that polarization fluctuations are responsible for other defining characteristics of bound electrons, such as their optical mass $\hbar/m\sim\langle{\bm{r}}\cdot{\bm{v}}\rangle$, part of orbital magnetization $M\sim\langle{\bm{r}}\times{\bm{v}}\rangle$, or shot noise $\langle{\bm{v}}\cdot{\bm{v}}\rangle$. As outlined in [38], the generalized sum rules can be systematically obtained from different time derivatives of the unequal time dipole-dipole correlator, or time-dependent QGT, $\mathcal{Q}_{\mu\nu}(t)=\langle\hat{r}_{\mu}(t)(1-\hat{P})\hat{r}_{\nu}\rangle$. Taking these derivatives amounts to convolute the geometric tensor with various powers of the energy differences at which the dipole transitions occur, generalizing the SWM sum rule one finds,
[…]
In the right-hand side band indices $n,m$ are summed over filled and empty states respectively, and the momentum dependence is implicit. In the left hand side the dissipative conductivity contains both the longitudinal and Hall components which pick up the symmetric and and antisymmetric parts of the geometric tensor. We introduce a low energy cutoff to remove any Fermi surface contribution. Assuming that the geometric contribution to the optical conductivity is dominated by a characteristic energy $\omega_{g}$, the optical mass may be approximated by $1/m\sim\omega_{g}\ell_{g}^{2}$, the shot noise by $\omega_{g}^{2}\ell_{g}^{2}$ and the dielectric constant of an insulator $\varepsilon=1+\chi$ with $\chi\sim\ell_{g}^{2}/\omega_{g}$. Note that this also highlights that moments with positive powers of energy differences $\omega^{\eta}$, $\eta>0$ become hard to estimate without accounting for the atomic contribution at high energies $\sim\omega_{a}^{\eta}\ell_{a}^{2}$. While this reasoning applies perfectly in a system with flat bands, such as in the Landau level problem, strictly speaking all different sum rules, or moments of ground state dipole fluctuations, should be computed by the knowledge of all states in the Hilbert space, both filled and empty, and this approximation may not be suited for highly dispersive bands and broad distributions of $\sigma(\omega)$. In this limit, we can instead keep track of these few moments as independent geometric scales, which can be used to accurately describe the response of materials beyond the semiclassical regime.
Other than estimations of geometric quantities, optical sum rules have been used to establish rigorous bounds on localization [39, 40, 41, 42], superfluid stiffness [43, 44, 45, 46], dielectric function [36, 47], and more recently, on the energy gap in topological insulators[48].
Geometric quantities also appear in density response functions at small $q$, as the density and position operators satisfy $\hat{\rho}_{q}\approx 1+iq_{\mu}\hat{r}_{\mu}$. This relation was recently employed to generalize sum rules to non-linear order using the density operator [49]. Furthermore, it provides a pathway to extract the QGT from the static structure factor [50, 51], which can then be utilized to establish topological bounds [52]. While such sum rules and related bounds are excellent for building intuition, they are challenging to measure or verify experimentally. Although some progress has been made [53, 54, 55], there is increasing evidence that extracting the quantum metric via sum rules in principle requires a sum over all frequencies [32].
IV.2 Transport coefficients
The quantum geometry of the ground state is an essential ingredient for the characterization of a bulk quantum phase. Nevertheless, its observational status remains rather complicated. These difficulties are intimately linked to its definition via the overlap between the projector $\hat{P}$ and the complement $1-\hat{P}$ (Eq. (3)), which means one should probe the system adiabatically, while simultaneously keeping track of all other energies. This somewhat paradoxical protocol is by all accounts hard to achieve but not impossible [56]. To give some orientation, in the following we will elucidate the main factors which enter into the integrand of a response function. However, to avoid an overly technical discussion, we will keep a loose definition of $\ell_{g}$, remembering that depending on the response, different and linearly independent geometric moments or even combinations of moments may contribute [57] which are usually tensorial objects and may differ from the definition in Eq. (7). We present in Table 1 the exact forms and references regarding these contributions.
Geometric quantities Observable Ingredients Definition
Berry curvature matrix elements $\Omega_{\mu\nu}^{nm}$ $\ell_{g}^{2}$ $(r_{\mu}^{nm}r_{\nu}^{mn}-r_{\nu}^{nm}r_{\mu}^{mn})$
Quantum metric matrix elements $g^{nm}_{\mu\nu}$ $\ell_{g}^{2}$ $(r_{\mu}^{nm}r_{\nu}^{mn}+r_{\nu}^{nm}r_{\mu}^{mn})$
band resolved Berry curvature $\Omega_{\mu\nu}^{n}$ $\ell_{g}^{2}$ $\sum_{m\neq n}\Omega_{\mu\nu}^{nm}$
band resolved quantum metric $g^{n}_{\mu\nu}$ $\ell_{g}^{2}$ $\sum_{m\neq n}g_{\mu\nu}^{nm}$
Landau Zener coupling $G_{\mu\nu}^{n}$ $\ell_{g}^{2}/\omega_{g}$ $\sum_{m\neq n}g_{\mu\nu}^{nm}/\omega_{nm}$
Quantum connection $Q^{nm}_{\mu\nu\lambda}$ $\ell_{g}^{3}$ $r^{nm}_{\mu}r^{mn}_{\nu}(r^{mm}_{\lambda}-r^{nn}_{\lambda}+i\partial_{\lambda}\operatorname{log}r_{mn}^{\mu})$
Phenomenon Observable Ingredients Definition Refs.
Anomalous Hall $\sigma_{\mu\nu}$ $\ell_{g}^{2}$ $\frac{e^{2}}{\hbar}\int_{\rm BZ}\sum_{n}f_{n}\Omega^{n}_{\mu\nu}$ [58, 59, 8, 60]
Nonlinear anom. Hall $\sigma^{(2)}_{\mu\mu\nu}$ $\ell_{\rm mr}\ell_{g}^{3}$ $\tau\frac{e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum_{n}f_{n}\partial_{\mu}\Omega^{n}_{\mu\nu}$ [61, 62, 63]
Static susceptibility $\chi_{\mu\nu}$ $\ell_{g}^{2}/\omega_{g}$ $\frac{e^{2}}{2\hbar}\int_{\rm BZ}\sum_{n}f_{n}G^{n}_{\mu\nu}$ [36]
Non-reciprocal conductivity $\sigma^{(2)}_{\mu\mu\nu}$ $\ell_{g}^{3}/\omega_{g}$ $\frac{e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum_{n}f_{n}\left(\partial_{\mu}G^{n}_{\mu\nu}-2\partial_{\nu}G^{n}_{\mu\mu}\right)$ [64, 65, 66, 67, 68, 69]
Optical transition rate $\Gamma_{\mu\nu}(\omega)$ $\ell_{g}^{2}/\omega_{\rm ext}$ $\pi\int_{\rm BZ}\sum^{\prime}_{nm}g^{mn}_{\mu\nu}\delta(\omega-\omega_{mn})$ [70, 71]
Current injection (linear pol.) $\sigma^{(2)}_{\mu\mu\lambda}(\omega)$ $\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$ $\tau\frac{2\pi e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}\partial_{\lambda}\omega_{mn}g_{\mu\nu}^{nm}\delta(\omega-\omega_{mn})$ [72, 73, 74, 75]
Current injection (circular pol.) $\sigma^{(2)}_{\mu\nu\lambda}(\omega)$ $\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$ $\tau\frac{2\pi e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}\partial_{\lambda}\omega_{mn}\Omega_{\mu\nu}^{nm}\delta(\omega-\omega_{mn})$ [76, 73, 74, 75]
Shift current $\sigma^{(2)}_{\mu\nu\lambda}(\omega)$ $\ell_{g}^{3}$ $\frac{i\pi e^{3}}{-2\hbar^{2}}\int_{\rm BZ}\sum^{\prime}_{nm}(Q^{nm}_{\mu\nu\lambda}-Q^{nm*}_{\nu\mu\lambda})\delta(\omega-\omega_{mn})$ [77, 78]
Spectral weight $n/m_{\lambda}$ $\ell_{g}^{2}\omega_{g}$ $(1/\hbar)\int_{\rm BZ}\sum^{*}_{nm}\omega_{mn}g^{nm}_{\lambda\lambda}$ [32]
Orbital magnetization $M_{\lambda}$ $\ell_{g}^{2}\omega_{g}$ $(e/c)\int_{\rm BZ}\sum^{*}_{nm}\omega_{mn}\Omega^{nm}_{\mu\nu}\epsilon_{\mu\nu\lambda}$ [79]
Localization tensor $g_{\lambda}$ $\ell_{g}^{2}$ $\int_{\rm BZ}\sum^{*}_{nm}g^{nm}_{\lambda\lambda}$ [57]
Chern number $\mathcal{C}_{\lambda}$ $\ell_{g}^{2}$ $\epsilon_{\mu\nu\lambda}\int_{\rm BZ}\sum^{*}_{nm}\Omega^{nm}_{\mu\nu}$ [80]
Table 1: Examples of geometric and mixed transport coefficients and some optical responses. Definitions: Occupation function $f_{n}$, difference of band energies $\hbar\omega_{nm}=\hbar\omega_{n}-\hbar\omega_{m}$, and the double sum over occupied and empty states is defined as $\sum^{*}_{nm}=\sum_{n,m\neq n}f_{n}(1-f_{m})$, while $\sum^{\prime}_{nm}=\sum_{n,m\neq n}(f_{n}-f_{m})$. To estimate the units of the response, divide by the volume on the unit cell and consider the units of its prefactors. The symbol $\epsilon_{\mu\nu\lambda}$ is the Levi-Civita totally antisymmetric tensor and $\delta_{mn}$ the Kronecker delta. Only the interband, geometric contributions.
Much progress has been made in diagnosing the most well-known part of the QGT - the Berry curvature ($\ell_{g}^{2}$), which can be recast as a single projection operation. As mentioned before, the Berry curvature leads to the intrinsic anomalous Hall effect, which describes the transverse, antisymmetric current component that arises in magnetic materials, which thus carry a nonzero momentum average of the Berry curvature. In magnetic insulators, the same mechanism gives rise to the quantum anomalous Hall effect [58, 59]. More recently, high harmonic generation and polarization-sensitive ARPES techniques were able to directly measure the momentum-resolved Berry curvature in SiO2 [81] and WSe2 [82]. Another route to measure the Berry curvature is to identify observables which couple to its momentum derivatives. For example, the nonlinear intrinsic anomalous Hall effect has been successfully expressed in terms of the quasiparticle lifetime times the momentum derivative of the Berry curvature ($\ell_{\rm mr}\ell_{g}^{3}$[61, 62, 63]. This derivative has become known as the Berry curvature dipole.
On the other hand, no direct measurements of the momentum-resolved quantum metric ($\ell_{g}^{2}$) or its momentum average have been reported so far. There have been suggestions that the metric might couple directly to the dc-conductivity [83, 84, 85] in certain fine-tuned settings, but no linear response observable is known which is generically determined solely by the quantum metric. Even though the quantum metric of the ground state seems rather elusive, the corresponding dipole matrix elements appear almost ubiquitously in response functions in the form of wavefunction overlaps. These matrix elements may have origin in quantum geometry or simply due to onsite atomic transitions, and are generally weighted by velocities or energies. The resulting observable is thus a mixed quantity which depends on quantum geometry and the band energies. To give one example, in a multi-orbital system, the nonlocal conductivity acquires a complicated dependence on several dispersive and geometric features [86]. Many cases of these mixed responses have been reported, but particularly successful have been examples where the interband quantum metric is normalized by the band gaps ($\ell_{g}^{2}/\omega_{g}$). Such a term accounts for the unavoidable mixing of bands in the presence of an adiabatic perturbation. This is highly analogous to the Landau-Zener effect [87] and is thus most appropriately viewed as the excitation of virtual interband transitions [88]. Landau Zener mixing can be identified in the low frequency linear capacitance of insulators, which has been employed to illuminate the role of the QGT for dielectric properties [36]. A similar effect can be observed in the nonlinear dc-conductivity of time-reversal breaking band structures [89, 75], leading to a novel nonlinear Hall effect and non-reciprocal longitudinal conductivity [64, 65]. Both contributions originate from the dipole of the Landau-Zener mixing ($\ell_{g}^{3}/\omega_{g}$[66, 67, 68, 69]. Recently, the third order conductivity has been derived for the study of altermagnets, which yields structurally analogous mixed response functions with Berry curvature and Landau-Zener terms as building blocks, albeit in more complicated combinations ($\ell_{g}^{4}/\omega_{g}$[90, 91].
Finally, one can search for responses which contain the band-resolved QGT in a given response function. Such is often possible in resonant optical responses. Notably, the interband matrix elements from optical transition rates (i. e. Fermi’s golden rule) yield the QGT, but evaluated with respect to two energies which are separated by the photon energy of the incident light (scale $\ell_{g}^{2}/\omega_{\rm ext}$[70, 71]. Furthermore, in materials breaking time-reversal symmetry, a nonreciprocal directional dichroism has been predicted which is proportional to the quantum metric times the quasiparticle velocity $(\ell_{g})$  [92, 93]. Several examples of resonant mixing have also been reported for the nonlinear optical conductivity [73, 74, 75]. Namely, there is a quantized circular photogalvanic effect in two-band systems originating from the Berry charges located at Weyl points ($\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$[76]. More generally, shift ($\ell_{g}^{3}$) and injection currents ($\ell_{\rm mr}\ell_{g}^{3}\omega_{\rm ext}$) have been connected to matrix elements of the QGT [72] and to geometric objects which go beyond the QGT, which has been termed Riemannian geometry  [94, 77, 95, 96] or multi-state geometry [97, 78].
Despite the tremendous progress in the last few years in characterizing geometric features in observables, it is important to stay parsimonious in applying this label. Matrix overlaps are ubiquitous in multi-orbital systems, and not every occurrence of a matrix element automatically signals that quantum geometry (dipole fluctuations from orbital frustration and delocalization) is an important ingredient towards understanding a particular response. We should be aware of the fact that the momentum structure of the matrix overlaps leads in many cases only to quantitative modifications, without changing the overall character of the response. The examples listed above show that it is nonetheless possible to distill and isolate the role of quantum geometry, and we are confident that many more examples like these will be found. However, to reiterate, wave function overlaps are a necessary, but not a sufficient ingredient for quantum geometric effects.
A second important caveat is the fact that almost all of the response functions discussed above can additionally receive large contributions from extrinsic origin, like impurity scattering, phonons and boundary effects [98, 63, 99]. It is an ongoing challenge to disentangle the latter from the intrinsic phenomena which originate from the bulk wavefunctions, and this is only possible on a case-by-case basis.
IV.3 Correlated phases
The SWM sum rule and the localization tensor have been known for over three decades. More recently, the surge of interest in quantum geometry has been largely driven by the discovery of the magic angle in twisted bilayer graphene (TBG) [100, 101, 102, 103] and the rapid development of moiré materials [104]. These systems exhibit flat bands due to destructive interference of electronic paths over large unit cells, where the geometric length scale is typically comparable to the unit cell size. While moiré materials are expected to host significant quantum geometry, much of the literature on quantum geometric effects differs conceptually from the theme of this review.
Our focus has been on dipole fluctuations in the ground state, whereas in moiré systems, the interest lies in the role of such fluctuations in modifying interactions that drive many-body phases. This arises because quantum geometry derives from dipole fluctuations which smear the real-space charge distribution, causing even short-range interactions to develop spatial profiles that can lead to a completely different ground state. Practically, these effects manifest through form factors, $\langle u_{k}|u_{k+q}\rangle$, in projected interactions, $\sum_{q}V_{q}\bar{\rho}_{q}\bar{\rho}_{-q}$ where the projected density operator is $\bar{\rho}_{q}=\sum_{k}\langle u_{k}|u_{k+q}\rangle c^{\dagger}_{k+q}c^{\phantom{\dagger}}_{k}$. Numerous studies on TBG have already highlighted the importance of form factors within a Hartree-Fock framework. There has also been intense effort to interpret these form factors in terms of well-understood systems like Landau levels [105] and heavy fermions [106].
The finite-$q$ overlaps that appear in projected density are fundamentally different from infinitesimal-$q$ overlaps that appear in projected position operator and the quantum geometric tensor [107]. As a result, while it all too common to find quantum geometric effects in projected interactions from form-factors, it is incredibly difficult to identify an observable that isolates the quantum geometric tensor. On dimensional grounds, such an observable must incorporate an additional length scale to account for the extra factor of $q$ that relates the projected density and position operators, $\bar{\rho}_{q}\approx 1+iq_{\mu}\cdot\bar{r}_{\mu}$.
The $f$-sum rule, defined in Eq. (8) with $\eta=1$, is one such observable with the right dimensions [38]. However, as noted in the previous section, the sum rule carries energetic pre-factors that make it different from the quantum metric. This is where interactions can come into play and replace the pre-factors with an interaction energy scale. While this argument is purely based on dimensional grounds, it is the working principle in flat-band superconductors.
A flat-band superconductor with attractive Hubbard interactions, $V_{q}=-U$, and the uniform pairing condition [108] is described by the Hamiltonian $H=-U\sum_{q}\bar{\rho}_{q}\bar{\rho}_{-q}$ [109]. The low-energy $f$-sum rule is highly sensitive to the quantum geometry of the non-interacting bands [44, 45]. In gapped BCS superconductors at $T=0$, it is directly related to the superfluid stiffness [110], which, within the mean-field approximation [111], is given by $D_{s}\sim n_{\rm 2D}U\ell_{g}^{2}$ where $n_{\rm 2D}$ is the charge density and $\ell_{g}^{2}$ is given by the trace of the quantum metric ${\rm tr}[g]$. Since the superfluid stiffness determines the effective mass of Cooper pairs, the geometric scale renormalizes the Cooper pair mass as $\hbar^{2}/m_{\rm eff}\sim U\ell_{g}^{2}$ [112].
Other than the sum rule and superfluid stiffness, several other observables, including coherence length, have been shown to exhibit quantum geometric corrections in a Landau-Ginzburg theory [113]. Although this framework of mass renormalization has been extensively applied to TBG [114, 115, 116], it is not limited to attractive interactions or superconducting ground states. Similar effects appear in systems with repulsive interactions, where the ground state includes excitons [117, 22, 118, 119].
The insights from TBG are now being extended to transition metal dichalcogenide (TMD) heterostructures, which exhibit a wide range of phenomena, including superconductivity [120], fractional quantum Hall phases [121, 122, 123, 124], quantum anomalous Hall effects [125], generalized Wigner crystals [126, 127], and various other correlated phases [128, 129, 130, 131, 132]. While form factors in projected densities are crucial for accurately modeling interactions [133], a comprehensive discussion of this growing body of work is beyond the scope of this review. For broader coverage of the field, we refer readers to [134].
We note in passing that moiré materials are not the only correlated systems where quantum geometric effects are apparent. Excitons in TMDs have long been observed to deviate from the effective mass approximation [135], with some of these deviations attributed to quantum geometric contributions [136, 137, 138]. Lastly, it is also becoming increasingly evident that a correlated ground state is not necessary for quantum geometric effects. The inter-band component of the Coulomb interaction can itself conspire with the inter-band current operator to renormalize the effective mass of carriers [139, 140]. This renormalization manifests in the plasmon group velocity of metals, which can be measured experimentally [141].
V Discussion and Outlook
The importance of quantum interference and orbital mixing across distinct lattice sites cannot be overstated. These quantum effects introduce electron dynamics at scales comparable to (and sometimes larger than) the lattice constant, fundamentally modifying the physical properties and phases of quantum materials.
Quantum geometry provides the mathematical framework to systematically account for these position fluctuation effects, both through matrix elements in correlation functions and through form factors in projected electron-electron interactions. As we emphasize in this overview, these effects can be understood by the length and time scales which are imposed not by the band dispersion, but by the momentum dependence of the electron wavefunctions. Therefore, we advocate to revisit minimal models of quantum materials from this fresh perspective. We should start by estimating the magnitude of geometric effects in materials with exotic physical properties. As described here, these scales can be quickly estimated using the optical conductivity (see Fig.3) both in ab-initio computations and direct measurement, and complemented by various other observables organized in Table 1.
That said, many fundamental questions remain open, and the full implications of quantum geometry in solid state physics are in its infancy. Much of the focus on quantum geometric effects have been given to insulators and flat bands [108], when in fact these are universal effects that dictate the impact of having multiple orbitals. Metals may be the most natural setting where quantum geometric effects can manifest, reformulating the conventional expectations of the response and correlated instabilities of various materials. Let us here list several key questions and directions we hope to see developments in the future:
Better models.
Quantum geometry quantifies the effects of Hilbert space truncation in effective theories, providing a systematic framework for including corrections that arise from orbital mixing. This can lead to more accurate effective models that capture the relevant scales without requiring the full atomic basis. If we are presented with materials with many orbitals and, at first glance, they cannot be disentangled, how do we answer the question: which orbitals matter?
Materials discovery.
Understanding the quantum geometry of materials and how it emerges by specific lattice motifs can guide the search for materials with enhanced responses or novel quantum phases. Large geometric effects often signal the presence of exotic physics, so how can we utilize the vast computational capabilities of material science to espouse geometric insights? Can we guess which materials will have which order based not only on fermiology but also on geometry?
Low temperature phenomena.
A key open challenge is understanding how interactions combined with significant dipole fluctuations affect the dynamics of electrons around the Fermi surface in metals. Can these effects be quantified through electronic transport at the Fermi surface, such as in its quantum oscillations? How can interactions with energies substantially smaller than the geometric energy $U\ll\hbar\omega_{g}$ lead to the transfer of spectral weight to zero frequencies, enhancing charge transport, both in metals [141] and correlated condensates [108, 22]? What is the mechanism by which geometry “speeds up” charge carriers in quantum matter?
Interpretation of experiments
Quantum geometry can help interpret experimental data, particularly when different probes appear to measure different effective masses or when responses deviate from semiclassical expectations.
Unifying Landau level and band phenomena.
Landau levels have extraordinary properties and are known to stabilize many exotic quantum states such as fractional topological order, exciton condensates [142]. However, it comes at the expense of high magnetic fields. The geometric framework introduces a geometric frequency $\omega_{g}$ and a geometric length $\ell_{g}$ that mimic the cyclotron frequency and magnetic lengths, however these emerge from lattice interference and not from the external Lorentz force. With the recent advent of moiré heterostructures, many quantum Hall phenomena have been mimicked in the absence of a magnetic field, opening a myriad of possibilities for new quantum devices. However, moiré systems come with their own problem of scalability. The proof of concept has been done - can we now find scalable materials with exotic orders which can be incorporated in quantum devices?
Let us conclude with the question: is it necessary to include quantum geometry in the standard toolbox? As physicists, we often look for the simplest description that abstracts away irrelevant degrees of freedom. Quantum geometry quantifies this process by allowing us to define energy and length scales with origin in orbital mixing. In modeling any condensed matter system, we strive to find a middle ground between complexity and simplicity, a compromise that dictates the level of description we adopt. Indeed, there is no ambiguity in defining the many-body Hamiltonian at the UV scale - the kinetic energy depends on the momentum operator, and interactions depend on the density operator. It is only when the theory is projected onto a subset of states that quantum geometric features emerge, often leading to low-energy theories that defy naive semiclassical intuition, as seen in flat-band superconductors and fractionalized quantum Hall states [143]
Quantum geometry is thus a natural and essential byproduct of our reliance on effective theories. Tying back to the notion of dipole fluctuations, this perspective can be succinctly stated as such: If the ground state of a material has strong orbital mixing, it cannot be described simultaneously on the basis of a single local orbital and in flat space. We can either choose a multi-band picture in flat space or a single-band picture in a curved geometry.